OpenStudy (anonymous):

HELP FOR A MEDAL , NOT SURE IF IM CORRECT :/ To open these doors, you must speak three functions in standard form. One function, f(x), with two real rational solutions. f(x)=x2-4 One function, g(x), with two real irrational solutions. g(x)= x2-2 One function, h(x), with two complex solutions. h(x)= -3x2+11i Create these three functions and explain to Professor McMerlock how you know these functions meet each condition. Remember, he is a Professor so use complete sentences. (Hint: Make sure that the b is even on g(x).)

OpenStudy (math&ing001):

Only few things you need to look at: 1. Your 'b' should be even on g(x). Standard form looks like this ax^2+bx+c your b=0 which is not an even number. 2. The h(x) function that you gave doesn't have a solution. It should look like h(x)=ax^2+b with 'a' and 'b' being positive.

OpenStudy (agent0smith):

@math&ing001 0 is an even number, as it is exactly divisible by 2.

OpenStudy (anonymous):

omg ok do you think you can help me out? :/ im really confused and i need to turn this in tomorrow

OpenStudy (agent0smith):

One function, f(x), with two real rational solutions. f(x)=x^2-4 that's correct, two solutions are x=2 and x=-2 One function, g(x), with two real irrational solutions. g(x)= x^2-2 correct, solutions are x= sqrt 2 and x = -sqrt 2 One function, h(x), with two complex solutions. h(x)= -3x2+11i This doesn't have real coefficients... just make something that never touches the x axis, like f(x) = x^2 + 1

OpenStudy (anonymous):

Ok thank you! & can you explain to me how to convert the first equation into the general, vertex form ?

OpenStudy (agent0smith):

The first is in vertex form. You could write it as \[\large f(x) = (x-0)^2 - 4\]but that's the same thing. All three are in vertex form.

OpenStudy (math&ing001):

@agent0smith yeah it's getting late I should probably get some sleep @omgkelley sry your second solution is correct and get the third function agent0 gave you. Good luck

OpenStudy (agent0smith):

You were prob just thinking of positive numbers (it's neither pos or neg.)

OpenStudy (anonymous):

ok thanks guys! :)